shinedivneonebytime:

A builder was building a fence. In the morning, he worked for 2/5 of an hour . in the afternoon he worked for 9/10 of an hour How many times as long did he work in the afternoon than in the morning

shinedivneonebytime:

im still confuse d

carmelle:

1 hour = 60 mins 60/5 = 12 12 + 12 = ?

shinedivneonebytime:

24

carmelle:

So that means the builder worked for 24 minutes in the morning

shinedivneonebytime:

im stil confused

shinedivneonebytime:

how am i supposed to do this

carmelle:

dividing

shinedivneonebytime:

can some one else help me

shinedivneonebytime:

in a diffrent way

justus:

Hello there. First, 9/10=2/5 so we divide both sides by 9/10, so would be 2/5÷9/10 = 4/9

shinedivneonebytime:

um im still confuesd why would we

shinedivneonebytime:

also i put it in and its coming up as wrong i put 4/9

justus:

In the morning, the builder worked 4/9 times as long as he worked in the afternoon.

shinedivneonebytime:

why would we divide both sides by 9/10

shinedivneonebytime:

just so confusing

jimthompson5910:

Consider a similar problem: In the morning he worked for 4 hours. In the afternoon he worked for 12 hours. How many times as long did he work in the afternoon than in the morning?

jimthompson5910:

I realize now that 12 hours in the afternoon isn't realistic, but let's ignore that bit for a moment.

shinedivneonebytime:

wait wait im confused why are you giving me a difffrent problem

jimthompson5910:

I'm using whole numbers instead of fractions, since fractions are often a pain to deal with.

shinedivneonebytime:

ok

shinedivneonebytime:

so how do answer the question im giving you

jimthompson5910:

as @carmelle was writing out 2/5 of an hour = (2/5)*60 = 24 minutes 9/10 of an hour = (9/10)*60 = 54 minutes He worked 24 minutes in the morning and 54 minutes in the afternoon How many times larger is 54 compared to 24?

shinedivneonebytime:

what do you mean by how much larger is 54 compared to 24?

jimthompson5910:

24 * ____ = 54 what number goes in the blank space?

jimthompson5910:

In other words, what is x in the equation 24x = 54 ?

shinedivneonebytime:

um 2.25

shinedivneonebytime:

?? hello

jimthompson5910:

correct

jimthompson5910:

so he worked 2.25 times longer in the afternoon compared to the morning

jimthompson5910:

as a fraction 2.25 = 9/4

shinedivneonebytime:

it says incorrect

shinedivneonebytime:

but how

jimthompson5910:

Did you type in 2.25? or 9/4?

shinedivneonebytime:

9/4

shinedivneonebytime:

and i tried typing 2.25 and it says wrong also

jimthompson5910:

Try 2.25 maybe. It's possible that your teacher wants a mixed number If so, then 9/4 = (8+1)/4 9/4 = 8/4 + 1/4 9/4 = 2 + 1/4 9/4 = 2 & 1/4

jimthompson5910:

0.25 = 1/4 = one quarter 2.25 = 2 and a quarter = 2 & 1/4

shinedivneonebytime:

thank you i have one more question

jimthompson5910:

go ahead

shinedivneonebytime:

A zookeeper is 6 and 1/4 feet tall A young giraffe in his care is 9 and 3/8 tallHow many times as tall as the zookeeper is the giraffe? ​

shinedivneonebytime:

i need help with this also

shinedivneonebytime:

helo

shinedivneonebytime:

hello

shinedivneonebytime:

heloooooo

shinedivneonebytime:

hello

carmelle:

hello

shinedivneonebytime:

hi

shinedivneonebytime:

can you help me

jimthompson5910:

sorry I was away, but I'll let @carmelle answer

carmelle:

she posted the question. I think someone is helping her

carmelle:

she/he